### v1

@@ 1.3 Organization
In Section 2 we present preliminaries necessary to prove our main results. In Section 3 we give a precise definition of memory capacity in the next-token prediction setting and prove the upper bound Theorem 1. In Section 4 we define the one-layer decoder-only transformer model. In Section 5 we prove the rank results. In Section 6 we prove the memory capacity results. In Section 7 we provide a numerical experiment demonstrating that not only is the memory capacity on the same order as the number of parameters, but the transformer can already be trained to the entropy lower bound in this regime. In Section 8, we conclude.
@@ 2 Preliminaries
Throughout the paper we use the following notation: $[n]$ denotes $\{1,\ldots,n\}$, $\Delta^{n-1}$ denotes the unit simplex in $\mathbb{R}^{n}$, $\mathrm{ri}\leavevmode\nobreak\ S$ denotes the relative interior of $S$, $\binom{A}{n}$ denotes $\{B\subset A\mid|B|=n\}$, vec denotes the column-wise vectorize operation, $e_{k}\in\mathbb{R}^{n}$ denotes the $k$th coordinate vector, $\mathbbm{1}_{n}$ denotes the vector of ones in $\mathbb{R}^{n}$, $I\in\mathbb{R}^{n\times n}$ denotes the identity matrix, $\delta$ denotes the Kronecker delta function, $a^{(k)}$ indicates that the exponent $k$ is applied to the vector $a$ element-wise, $\mathfrak{S}_{n}$ denotes the symmetric group of degree $n$, $\exp$ denotes the exponential function, and $\phi$ denotes the softmax function,
@@
We also use the following conventions for matrices. If $a_{i}\in\mathbb{R}\leavevmode\nobreak\ \forall i\in[m]$, then $[a_{i}]_{i\in[m]}$ denotes $(a_{1},\ldots,a_{m})\in\mathbb{R}^{m}$. If $a_{j}\in\mathbb{R}^{m}\leavevmode\nobreak\ \forall j\in[n]$, then $[a_{j}]_{j\in[n]}$ denotes $[a_{1}|\cdots|a_{n}]\in\mathbb{R}^{m\times n}$. Going the other way, if $a\in\mathbb{R}^{m}$, then $a_{i}$ denotes element $i$ of $a$, and if $A\in\mathbb{R}^{m\times n}$, then $a_{i,j}$ denotes element $(i,j)$ of $A$. We will also index vectors and matrices using brackets. If $a\in\mathbb{R}^{m}$ and $\alpha\in[m]^{k}$, then $a[\alpha]$ denotes $[a_{\alpha_{i}}]_{i\in[k]}$ and $\mathrm{len}(a)$ denotes $m$. If $A\in\mathbb{R}^{m\times n}$, $\alpha\in[m]^{k}$, and $\beta\in[n]^{\ell}$, then $A[\alpha,\beta]$ denotes $[[a_{\alpha_{i},\beta_{j}}]_{i\in[k]}]_{j\in[\ell]}$. In this context, a colon denotes $(1,2,\ldots)$ and $i:j$ denotes $(i,\ldots,j)$.
In the realm of real analytic functions, we use Taylor’s theorem as as well as the fact that the zero set of a nontrivial real analytic function is measure zero and closed (Gunning and Rossi, 1965, Corollary 10). We say a property holds “generically” if the set of exceptions lies on the zero set of a nontrivial real analytic function. Note that $\phi$ is real analytic since the numerator and denominator are real analytic with the denominator being nowhere zero. Moreover, the entire self-attention sub-layer of the transformer model, which will be described in detail below, is real analytic since it is a composition of real analytic functions. Furthermore, if $\psi$ is real analytic at a point, then the entire transformer parameter mapping is real analytic at a point. Similarly, if $\psi$ is continuously differentiable, then the entire transformer parameter mapping is continuously differentiable, in which case we can apply Sard’s theorem (Sard, 1942).
@@
In the realm of linear algebra, we use the notion of Kruskal rank (Kruskal, 1977), as well as the Cauchy-Binet formula (Gantmacher, 1960, Sec. I.2.4) and the Leibniz determinant formula (Axler, 2015, Def. 10.33). In the realm of number theory, we use that $\exp(1)$ is transcendental (Hermite, 1873).
3 Next-token prediction
We precisely define the memory capacity of a model in the next-token prediction setting in the following way.
Definition 1.
Let $k,\omega,T\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Let $\phi$ be the softmax function. Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is the largest $n\in\mathbb{N}$ such that for all $\tau_{1},\ldots,\tau_{n}\in[T]$ and distinct $\alpha^{1}\in[\omega]^{\tau_{1}},\ldots,\alpha^{n}\in[\omega]^{\tau_{n}}$,
$\{[(\phi\circ h_{\theta})(\alpha^{1})|\cdots|(\phi\circ h_{\theta})(\alpha^{n})]\mid\theta\in\mathbb{R}^{k}\}=(\mathrm{ri}\Delta^{\omega-1})^{n}.$
We call $[\omega]$ the “vocabulary” and its elements “tokens.” The tokens may represent letters, words, or something in between, as in Byte Pair Encoding (Gage, 1994). Given a sequence of tokens $\alpha\in[\omega]^{\tau}$, we think of $(\phi\circ h_{\theta})(\alpha)\in\mathrm{ri}\Delta^{\omega-1}$ as the model’s prediction of the conditional probability distribution of the token to follow $\alpha$. If $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$, then we have the following upper bound on the memory capacity.
Lemma 1.
Let $k,\omega,T\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$. Let $\phi$ be the softmax function. Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is less than or equal to $k/(\omega-1)$.
@@ 3 Next-token prediction
Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{T}[\omega]^{t}$ be distinct. Define $F$ as the mapping $\theta\mapsto[(\phi\circ h_{\theta})(\alpha^{1})|\cdots|(\phi\circ h_{\theta})(\alpha^{n})]$. $F$ is continuously differentiable, so we can apply Sard’s theorem (Sard, 1942). Note that $\mathrm{ri}\Delta^{\omega-1}$ is an $(\omega-1)$-dimensional manifold. Thus, if $k<n(\omega-1)$, then the image of $F$ has measure zero, hence $F$ is not surjective. ∎
In practice, we are generally given $N$ sequences $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$, and $\theta$ is trained to minimize the cross-entropy loss,
$\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})\coloneqq-\sum_{i=1}^{N}\sum_{t=1}^{T}\log\left(\phi\left(h_{\theta}\left(\beta_{1}^{i},\ldots,\beta_{t}^{i}\right)\right)_{\beta_{t+1}^{i}}\right).$
Enumerate the multiset $\{(\beta_{1}^{i},\ldots,\beta_{t}^{i})\mid i\in[N],t\in[T]\}$ as $\{(\alpha^{1},c_{1}),\ldots,(\alpha^{n},c_{n})\}$. Let $\tau_{i}$ denote the length of $\alpha^{i}$ for each $i\in[n]$ and define
$\hat{\pi}_{\gamma,i}\coloneqq\frac{|\{j\in[N]\mid(\beta_{1}^{j},\ldots,\beta_{\tau_{i}}^{j})=\alpha^{i},\beta_{\tau_{i}+1}^{j}=\gamma\}|}{c_{i}}\leavevmode\nobreak\ \forall(\gamma,i)\in[\omega]\times[n].$ (1)
For each $i\in[n]$, define $\hat{\pi}_{i}=[\hat{\pi}_{\gamma,i}]_{\gamma\in[\omega]}$. We call the $\alpha^{1},\ldots,\alpha^{n}$ the “contexts” of $\beta^{1},\ldots,\beta^{N}$. Then $c_{i}$ is the number of times that context $\alpha^{i}$ appears and $\hat{\pi}_{i}$ is the empirical distribution of the next-token conditioned on $\alpha^{i}$. Define the entropy
$\mathcal{H}(\beta^{1},\ldots,\beta^{N})\coloneqq-\sum_{i=1}^{n}\sum_{\gamma\leavevmode\nobreak\ :\leavevmode\nobreak\ \hat{\pi}_{\gamma,i}\neq 0}c_{i}\hat{\pi}_{\gamma,i}\log\left(\hat{\pi}_{\gamma,i}\right).$
Then the following lemma follows from Gibb’s inequality.
Lemma 2.
Let $k,\omega,T,N\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Let $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$. Then,
$\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})\geq\mathcal{H}(\beta^{1},\ldots,\beta^{N})\leavevmode\nobreak\ \forall\theta\in\mathbb{R}^{k}.$
Moreover, if $\{[(\phi\circ h_{\theta})(\alpha^{1})|\cdots|(\phi\circ h_{\theta})(\alpha^{n})]\mid\theta\in\mathbb{R}^{k}\}=(\mathrm{ri}\Delta^{\omega-1})^{n}$, then
$\inf_{\theta\in\mathbb{R}^{k}}\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})=\mathcal{H}(\beta^{1},\ldots,\beta^{N}).$
Applying Lemma 2 to a memory capacity lower bound results in an “empirical memory capacity” lower bound, which can be compared to the following empirical memory capacity upper bound. Note that we have to further assume that $\theta\mapsto h_{\theta}(\alpha)$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$. As an example, this is satisfied by the transformer model described in the next section for activations such as tanh, arctan, and GELU.
Theorem 1.
Let $k,\omega\in\mathbb{N}$. Set $n^{\prime}=\lfloor k/(\omega-1)\rfloor+1$ and $T=\lceil\log(n^{\prime})/\log(\omega)\rceil$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$. Then there exists $N\geq n^{\prime}$ and $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$ such that the number of contexts is less than or equal to $(2+1/(\omega-1))k/(\omega-1)+1$ and
$\inf_{\theta\in\mathbb{R}^{k}}\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})>\mathcal{H}(\beta^{1},\ldots,\beta^{N}).$

### v2

@@ 1.3 Organization
In Section 2 we present notations, conventions, and prerequisites. In Section 3 we carefully delineate the assumptions made to start talking about next-token prediction. Section 3 serves to ground the definition of next-token prediction capacity, but the reader that is already willing to accept the definition can safely skip to Section 4. In Section 4 we define next-token prediction capacity and prove the upper bound Theorem 4.5. In Section 5 we define the one-layer multi-head decoder-only transformer model. In Section 6 we prove the injectivity result, the rank results, and the next-token prediction lower bound. In Section 7 we present a numerical experiment demonstrating that not only is the memory capacity on the same order as the number of parameters divided by the vocabulary size, but the transformer can already be trained to the entropy lower bound in this regime. In Section 8 we introduce the token-averaged FNN model and prove a next-token prediction capacity lower bound for it. In Section 9, we conclude.
@@ 2 Preliminaries
[…]
We will use the following notation: $[n]$ denotes $\{1,\ldots,n\}$, $\Delta^{n-1}=\{x\in\mathbb{R}_{+}^{n}\mid\|x\|_{1}=1\}$ denotes the unit simplex in $\mathbb{R}^{n}$, $\mathrm{ri}\Delta^{n-1}=\{x\in\mathbb{R}_{++}^{n}\mid\|x\|_{1}=1\}$ denotes the relative interior of $\Delta^{n-1}$, $\binom{A}{n}$ denotes $\{B\subset A\mid|B|=n\}$, vec denotes the column-wise vectorize operation, $e_{k}\in\mathbb{R}^{n}$ denotes the $k$th coordinate vector, $\mathbbm{1}_{n}$ denotes the vector of ones in $\mathbb{R}^{n}$, $I\in\mathbb{R}^{n\times n}$ denotes the identity matrix, $\delta$ denotes the Kronecker delta function, $a^{(k)}$ indicates that the exponent $k$ is applied to the vector $a$ element-wise, $\mathfrak{S}_{n}$ denotes the symmetric group of degree $n$, $\exp$ denotes the exponential function, and $\phi$ denotes the softmax function,
@@
We also use the following conventions for matrices. If $a_{i}\in\mathbb{R}\leavevmode\nobreak\ \forall i\in[m]$, then $[a_{i}]_{i\in[m]}$ denotes $(a_{1},\ldots,a_{m})\in\mathbb{R}^{m}$. If $a_{j}\in\mathbb{R}^{m}\leavevmode\nobreak\ \forall j\in[n]$, then $[a_{j}]_{j\in[n]}$ denotes $[a_{1}|\cdots|a_{n}]\in\mathbb{R}^{m\times n}$. Going the other way, if $a\in\mathbb{R}^{m}$, then $a_{i}$ denotes element $i$ of $a$, and if $A\in\mathbb{R}^{m\times n}$, then $a_{i,j}$ denotes element $(i,j)$ of $A$. We will also index vectors and matrices using brackets. If $a\in\mathbb{R}^{m}$ and $\alpha\in[m]^{k}$, then $a[\alpha]$ denotes $[a_{\alpha_{i}}]_{i\in[k]}$. If $A\in\mathbb{R}^{m\times n}$, $\alpha\in[m]^{k}$, and $\beta\in[n]^{\ell}$, then $A[\alpha,\beta]$ denotes $[[a_{\alpha_{i},\beta_{j}}]_{i\in[k]}]_{j\in[\ell]}$. In this context, a colon denotes $(1,2,\ldots)$, $i:j$ denotes $(i,\ldots,j)$, and $-1$ denotes the last index.
In the realm of real analytic functions, we use Taylor’s theorem as as well as the fact that the zero set of a nontrivial real analytic function is measure zero and closed (Gunning and Rossi, 1965, Corollary 10). We say a property holds “generically” if the set of exceptions lies on the zero set of a nontrivial real analytic function. Note that $\phi$ is real analytic since the numerator and denominator are real analytic with the denominator being nowhere zero. Moreover, the entire self-attention sub-layer of the transformer model, which will be described in detail below, is real analytic since it is a composition of real analytic functions. Furthermore, if $\psi$ is real analytic at a point, then the entire transformer parameter mapping is real analytic at a point. Similarly, if $\psi$ is continuously differentiable, then the entire transformer parameter mapping is continuously differentiable, in which case we can apply Sard’s theorem (Sard, 1942), which states that the set of critical values has measure zero. A critical point is a point where the differential is not surjective and a critical value is a value with level set (preimage/fiber) containing at least one critical point.
@@
In the realm of linear algebra, we use the notion of Kruskal rank (Kruskal, 1977), as well as the Cauchy-Binet formula (Gantmacher, 1960, Sec. I.2.4) and the Leibniz determinant formula (Axler, 2015, Def. 10.33). In the realm of number theory, we use that $\exp(1)$ is transcendental (Hermite, 1873). Finally, we also use basic concepts from probability theory to which we refer the reader to Billingsley (1995) as a reference.
3 Probabilistic language space
We start by abstracting notions about how language works into mathematical axioms. Specifically, given a sequence of words, we assume that the next word is governed by a probability distribution. Of course, this is not how language really works—we respond to another person’s body language in conversation and the author writes about flowers they see on the side of the road—but it is a useful abstraction for certain tasks such as summarizing material online. We will axiomatize the concept of a probabilistic language space in three different ways and show that they are equivalent. Note that we assume neither stationarity nor ergodicity. Refer to Section 1.2 for the context of this section’s results within the related literature on the probability of language, tracing back to the seminal papers of Markov and Shannon. Also, note that, while this section serves to ground the concept of next-token prediction capacity, the reader who wishes to take the definition of next-token prediction capacity at face value can safely skip to the next section.
Definition 3.1.
A probabilistic language space of the first kind is a triple $(\mathcal{X},A,p)$ that satisfies the following axioms:
  1. 1.
    $\mathcal{X}$ is a finite nonempty set;
  2. 2.
    $\mathcal{X}^{0}\subset A\subset\cup_{t=0}^{\infty}\mathcal{X}^{t}$;
  3. 3.
    $p:\mathcal{X}\times A\to[0,1]$;
  4. 4.
    $\forall x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$, $x\in A\iff p(x_{t}\mid x_{1}\ldots,x_{t-1})\neq 0$;
  5. 5.
    $\forall x\in A$, $\sum_{y\in\mathcal{X}}p(y\mid x)=1$.
Definition 3.2.
A probabilistic language space of the second kind is a couple $(\mathcal{X},p)$ that satisfies the following axioms:
  1. 1.
    $\mathcal{X}$ is a finite nonempty set;
  2. 2.
    $p:\cup_{t=1}^{\infty}\mathcal{X}^{t}\to[0,1]$;
  3. 3.
    $\sum_{x\in\mathcal{X}}p(x)=1$;
  4. 4.
    $\forall x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$, $\sum_{y\in\mathcal{X}}p(x,y)=p(x)$.
Definition 3.3.
A probabilistic language space of the third kind is a couple $(\mathcal{X},P)$ that satisfies the following axioms:
  1. 1.
    $\mathcal{X}$ is a finite nonempty set;
  2. 2.
    $P$ is a probability measure on the cylindrical $\sigma$-algebra of $\mathcal{X}^{\mathbb{N}}$;
  3. 3.
    $\sum_{x\in\mathcal{X}}P(\mathcal{C}(\{x\}))=1$;
  4. 4.
    $\forall x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$, $\sum_{y\in\mathcal{X}}P(\mathcal{C}(\{(x,y)\}))=P(\mathcal{C}(\{x\}))$;
where, for all $t\in\mathbb{N}$ and all $A\subset\mathcal{X}^{t}$, $\mathcal{C}(A)\coloneqq\{x\in\mathcal{X}^{\mathbb{N}}\mid(x_{1},\ldots,x_{t})\in A\}$.
Axiom 5 of the first definition says that $p(\cdot\mid x)$ is a conditional probability mass function whenever it is defined. Axiom 4 says that $p(\cdot\mid x)$ is defined for all $x$ except those which have zero probability by the chain rule. Definition 3.2 essentially applies the chain rule to the conditional probability mass functions from Definition 3.1. Note that for all $t\geq 2$, $\sum_{x\in\mathcal{X}^{t}}p(x)=\sum_{y\in\mathcal{X}^{t-1}}\sum_{z\in\mathcal{X}}p(y,z)=\sum_{y\in\mathcal{X}^{t-1}}p(y)$, so, by induction, $\sum_{x\in\mathcal{X}^{t}}p(x)=1$ for all $t\in\mathbb{N}$. In other words, $p$ is a probability mass function on $\mathcal{X}^{t}$ for every $t\in\mathbb{N}$, but also is compatible with being decomposed into the conditional probability mass functions of Definition 3.1.
The third definition requires greater explanation. Subsets of $\mathcal{X}^{\mathbb{N}}$ of the form $\mathcal{C}(A)$ for some $A\subset\mathcal{X}^{t}$ are called cylinders of rank $t$. Subsets of $\mathcal{X}^{\mathbb{N}}$ of the form $\mathcal{C}(\{x\})$ for some $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$ are called thin cylinders. Let $\mathscr{C}_{0}$ be the class of cylinders. Then $\mathscr{C}_{0}$ is an algebra, as shown after Eq. (2.17) in Billingsley (1995). The $\sigma$-algebra generated by it, $\mathscr{C}$, is called the cylindrical $\sigma$-algebra of $\mathcal{X}^{\mathbb{N}}$. Axioms 3 and 4 of Definition 3.3 ensure that $P$ can be reduced to a $p$ as in Definition 3.2.
It is quite easy to show that the first two definitions are equivalent, but distinguishing them so formally is useful because it makes the connection between the last two definitions more clear. Normally, $p$ is extended to $P$ without proving that there is actually a bijection between the set of $p$ such that $(\mathcal{X},p)$ is a probabilistic language space of the second kind and the set of $P$ such that $(\mathcal{X},P)$ is a probabilistic language space of the third kind. We will rigorously prove this after proving the equivalence of the first two definitions.
Proposition 3.4.
Let $\mathcal{X}$ be a finite nonempty set. Define $\mathbb{L}_{1}$ as the set of all $p:\mathcal{X}\times A\to[0,1]$ such that $(\mathcal{X},A,p)$ is a probabilistic language space of the first kind. Define $\mathbb{L}_{2}$ as the set of all $p:\cup_{t=1}^{\infty}\mathcal{X}^{t}\to[0,1]$ such that $(\mathcal{X},p)$ is a probabilistic language space of the second kind. Then there is a bijection $\Phi_{12}:\mathbb{L}_{1}\to\mathbb{L}_{2}$ such that for all $p:\mathcal{X}\times A\to[0,1]$ in $\mathbb{L}_{1}$,
$\Phi_{12}(p)(x,y)=p(y\mid x)\Phi_{12}(p)(x)\leavevmode\nobreak\ \forall(y,x)\in\mathcal{X}\times A.$
@@ 3 Probabilistic language space
Define $\Phi_{12}:\mathbb{L}_{1}\to\mathbb{L}_{2}$ in the following way. Let $p:\mathcal{X}\times A\to[0,1]$ be in $\mathbb{L}_{1}$. Extend $p$ to all of $\mathcal{X}\times\cup_{t=0}^{\infty}\mathcal{X}^{t}$ by setting $p(y\mid x)=0$ for all $(y,x)\in\mathcal{X}\times A^{\mathsf{c}}$. Then, for all $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$, define $q(x)=\prod_{t=1}^{|x|}p(x_{t}\mid x_{1},\ldots,x_{t-1})$. Then $(\mathcal{X},q)$ satisfies the first three axioms of Definition 3.2 by construction. To see that it satisfies the fourth axiom as well, let $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$ and observe
$\sum_{y\in\mathcal{X}}q(x,y)=q(x)\sum_{y\in\mathcal{X}}p(y\mid x)=q(x).$
Thus, $q\in\mathbb{L}_{2}$. Define $\Phi_{12}(p)=q$.
We will show that $\Phi_{12}$ is a bijection by constructing its inverse. Define $\Phi_{21}:\mathbb{L}_{2}\to\mathbb{L}_{1}$ in the following way. Let $q\in\mathbb{L}_{2}$. Extend $q$ to all of $\cup_{t=0}^{\infty}\mathcal{X}^{t}$ by setting $q(\leavevmode\nobreak\ )=1$. Define $A=\mathcal{X}^{0}\cup\{x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}\mid q(x)\neq 0\}$ and $p:\mathcal{X}\times A\to[0,1]:(y,x)\mapsto q(x,y)/q(x)$. Then $p$ is well-defined and satisfies the first four axioms of Definition 3.1 by construction. To see that it satisfies the fifth axiom, we will use induction. First, $\sum_{y\in\mathcal{X}}p(y\mid\leavevmode\nobreak\ )=\sum_{y\in\mathcal{X}}q(y)=1$. Now let $t\in\mathbb{N}$ and suppose $\sum_{y\in\mathcal{X}}p(y\mid x)=1$ for all $x\in A$ such that $|x|<t$. Then let $x\in A$ such that $|x|=t$ and observe
$\sum_{y\in\mathcal{X}}p(y\mid x)=\sum_{y\in\mathcal{X}}\frac{q(x,y)}{q(x)}=1.$
Thus, $p\in\mathbb{L}_{1}$. Define $\Phi_{21}(q)=p$. Clearly, $\Phi_{21}$ is the inverse of $\Phi_{12}$. ∎
Proposition 3.5.
Let $\mathcal{X}$ be a finite nonempty set. Define $\mathbb{L}_{2}$ as the set of all $p:\cup_{t=1}^{\infty}\mathcal{X}^{t}\to[0,1]$ such that $(\mathcal{X},p)$ is a probabilistic language space of the second kind. Define $\mathbb{L}_{3}$ as the set of all probability measures on $(\mathcal{X}^{\mathbb{N}},\mathscr{C})$ such that $(\mathcal{X},P)$ is a probabilistic language space of the third kind. Define $\Phi_{32}:\mathbb{L}_{3}\to\mathbb{L}_{2}:P\mapsto P(\mathcal{C}(\{\cdot\}))$. Then $\Phi_{32}$ is a bijection.